UT 20160424 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson,
Observer AO: DMiller, GTaylor (both LBTO), EPinna, APuglisi (remote from Italy upn request)
Telescope Operator: SAllenson
LBTO Support: TSargent

Plan for the night

As it is very windy we may not open, thus:

- AFC- and quite a few other closed-dome tests

...and if we are gonna to open:

- AFC-tests+ stability of system

Summary

- Night basically lost due to adverse observing conditions and technical issues

Details

All times are UT
Files are luci1.20160424.NNNN.fits

02:00 Sunset, mostly clear, T = +4deg, Hum = 60%, strong winds ~16-20m/s + higher gusts,
need to stay closed for the moment being

04:30 Wind has calmed down a bit, we are going to open, but have restrictions on sky due to Wind (go for NE),
Steve checks pointing and collimates, DIMM reports 1.5-2"

AFC + stability of system

DT prepared a script which does AFC in ima mode with LUCI1 and spectroscopy on LUCI2 (both AO).
Imaging is done in zJHK, spectroscopy with G200+zJspec centered at 1.07mue ([O II]. The latter doesn't
require NCPA (not available for LUCI2 yet). Its a bino-script and will be tried once, if it fails,
we will stop and do imaging with LUCI1 only. Poor observing conditions, thus this test is
prefectly appropriate.

04:50 Preset to 3c294_bino, worked, RunAO, left side failed (timeout bayside Y), right side completed (ESM only),
left side completed, LUCI1: Bin=3, R=13.0; LUCI2: Bin=2, R=12.5, though same star,
need to resent preset (due to failed RunAO on the left side), OK, RunAO, left side failed (skip frame while doing autogain),
resent preset again, OK, RunAO, left side failed once more (skip frame), seeing too bad, too much forces on secondary.
No way to do AO-stuff!! OK, switch to LUCI2 NIR-spectroscopy

05:45 Wind again above limits, need to close, Cirrus coming in too

Dave did some closed-dome tests (logs will come in later)

08:30 Wind below limits, we open again, both LUCIs are authorized, Steve has problems with the (co)pointing,
problems with the collimation on the right side, restart GCS, problem still persists, flattened shell, nope

09:36 Gosh, had to close again, humidity above limits!

Steve and Doug went up to inspect the right side, all looks OK

10:15 Humidity has dropped, Steve opens again, since we may use LUCI2 later this run,
we wil first troubleshoot to see whats wrong, checks pointing...all good, who understand this,
DIMM seeing varies wildly between 1.5 and 3", no way for doing AO

N30/G210 test-spectra LUCI2

10:44 Preset to MRK290, OK, #0028 throughslit, #0029 slitview, #0030 verification, OK,
N30/G210: J@1.25mue, #0031-0032, J@1.65mue, #0033-0034, some nightsky emission lines in H-band

12:05 Steve is closing up

-- %USERSIG{JochenHeidt - 2016-04-24}%
Topic revision: r6 - 18 Oct 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback